A pump move warm water at 120F from a reservoir open to atmosphere. The pipe is 6" diameter cast and has a standard 90 degree elbow and a fully open gate valve. The pump draws water from a pipe submerged below the water surface. The pump is 10 feet above the reservoir surface, and there are 20 feet of pipe in the inlet section. If the flow rate is 1,650 gpm, what is the NPSHA at the pump inlet?
Answer #1

Given data: Cast iron pipe size, D = 6 = 0.5 ft. Pipe length, L = 20 ft. Flow rate, Q = 1650 gpm Solution: The equation of ND2 220.57 20.52 V = 1123 ft/min V = 18.72 ft/s v = kinematic viscosity = 6.06 * 10 ft/s R = (0.5*18.72)/6.06*10* = 1544765 Itho=29 18.622 hv = 2 * 32 hy = 5.4 ft Head loss due to 90° bend is hf2Fkoh,=0.7*5.4 = 3.78 Calculate vapor pressure head, he h